Community acquired methicillin resistant Staphylococcus aureus (CA-MRSA) is a cause of skin infections. A case-control studyCalculate the OR and 95% Confidence Interval for MRSA for cat owners compared to non-cat owners. Please show your work.
Answer #1

Assuming significance of 0.05 since it's not given in the question.

1) Not Exposed to children case - Relationship between CAT in household and CA-MRSA

Chi-Square statistic for this equals 0.547

Degrees of Freedom = (2-1)*(2-1) = 1

Chi-square critical stat at 1 degrees of freedom = 0.004

0.547>0.004, hence, rejecting null that there is no association between CAT in household and CA-MRSA. Hence, there is association between CAT in household and CA-MRSA.

2) Exposed to children case - Relationship between CAT in household and CA-MRSA

Chi-Square statistic for this equals 0.722

Degrees of Freedom = (2-1)*(2-1) = 1

Chi-square critical stat at 1 degrees of freedom = 0.004

0.722>0.004, hence, rejecting null that there is no association between CAT in household and CA-MRSA. Hence, there is association between CAT in household and CA-MRSA.